OpenStudy (anonymous):

Which of the following is equivalent to five and four-ninths divided by seven? seven-ninths one and two-sevenths five-sevenths five and four sixty-thirds

OpenStudy (anonymous):

49/9 divided by 7 = 49/9 d multilplied by 1/7 = 7/9 = a

OpenStudy (anonymous):

you can medal me if you like XD

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

thanks!!